# Input-Output Optics as a Causal Time Series Mapping: A Generative Machine Learning Solution

### v1

@@ I Introduction
Optics studies light and its interactions with matter. As depicted in Fig 1, the study comprises three main components: a quantum system, the incident pulse of light (optical input), and the system’s response (optical output) to the incident pulse. The system’s response usually varies depending on the input signal. The type of input signal can determine whether the dynamical system behaves linearly, non-linearly, or chaotically (a specific type of non-linear behavior characterized by extreme sensitivity) [1]. If the input is weak, we have what is called linear optics. If the input intensity increases, we have a non-linear response/output signal, and the output can be expressed as a perturbation series expansion of the input [2]. As the intensity of the input field increases further, the system exhibits even more pronounced non-linearities requiring non-pertrubative treatment [3], one of the most significant being High Harmonic Generation (HHG). This strong-field laser–matter interaction resulting in HHG has attracted a lot of attention, and its discovery has been recognized with the 2023 Nobel Prize in Physics. HHG has been extensively studied in various media, including gases, semiconductors, nano-structures, metal, strongly correlated materials, and solids [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. More recently, there is mounting interest in HHG from quantum spin systems due to their potential applications in probing spin dynamics and developing new laser sources in the THz regime [27, 28, 29, 30].
In ultrafast spectroscopy, HHG is used to create attosecond pulses [31], which allow for unmatched temporal resolution of the observation of electron dynamics in real-time. Also, recent analysis suggests that HHG enables single-atom computing [32] and improved spectroscopic characterization of mixtures [33]. Furthermore, coherent extreme ultraviolet and soft X-ray light generation is made possible by HHG, and is an essential technology for high-resolution imaging and nanolithography in the semiconductor manufacturing industry.
@@
The optical output (Fig 1) is obtained theoretically by solving the time-dependent Schrödinger equation. Solving this equation rapidly becomes prohibitively expensive, as finding its solution is exponentially difficult with respect to the size of the quantum system with which the input pulse interacts. The complexity increases further if the quantum system is undergoing a phase transition. HHG of a phase transition has also been observed experimentally [34]. Understanding phase transitions is important for various reasons [35, 36, 37, 38]. For example, in quantum computing, the transverse field Ising model is significant for quantum annealing, where phase transitions can influence the performance and efficiency of quantum algorithms [39]. The transverse Ising model is well known to exhibit a quantum phase transition and is analytically solvable [40]. However, an Ising model with mixed transverse and longitudinal magnetic fields can only be solvednumerically  [41]. In such regimes, small changes in system parameters can lead to large-scale changes in the system’s dynamics, requiring extremely precise calculations that are challenging to achieve by numerically solving the Schrödinger equation. Alternative methods are needed to accurately model the behavior of such systems.
@@
In this paper, we propose a novel data-driven methodology to model the input $\to$ output dynamics of optical systems using machine learning. We utilize Temporal Convolutional Networks (TCNs) to capture the intricate interactions in many-body quantum systems by formulating optics as an input-output time series problem. Additionally, we demonstrate how distinct physical regimes, including phase transitions, can be captured with machine learning models (see Fig. 1), tailored to the complexity of each regime.
@@
A conventional application of machine learning in time-series analysis is forecasting [51, 52]. A notable example is Ref. [53], where using a historical record (e.g., the training data shown in Fig. 2) for a time-domain HHG output $Y(t)$, a model was constructed to predict $Y(t)$ for future times (see Fig. 2). It is important to note that this modeling does not utilize information across different input pulses. Hence, each time the incident laser pulse is changed, a new model needs to be trained. In the present work, we depart from this forecasting approach by training a TCN to learn the entire input $\to$ output relationship, employing a dataset of pairs of input and output pulses.
Figure 2: Schematic picture of time series modelling using TCN/LSTM: Y(t) along the y-axis is the output (for example HHG spectra) and Time is along the x-axis (a.u represents atomic units). The blue portion depicts the historical data that serves as the training data for the model, the red portion depicts the test data or the unseen data for the model and the green curve depicts the model’s predicted data whose closeness to the test data is a measure of the performance of the model.
@@
The application of machine learning to HHG has been prolific. However, unlike our current work, the previous studies have not enforced causality. These include applications in harmonic generation microscopy [65], high harmonic spectroscopy in solids [66], predicting characteristics of harmonic generation in plasmas using particle-in-cell simulations [67], predicting HHG emission from molecules and inversely predicting molecular parameters from the corresponding HHG spectra [68], time series forecasting of HHG spectra [53], and predicting HHG from the spatially structured input fields [69].
@@
The rest of the paper is organized as follows: In Sec. II, we comprehensively formulate the problem under investigation and provide details of the Ising system under consideration and various cases associated with it. Section III is dedicated to presenting our results; we state details about the machine learning model architectures employed (Sec. III.1), evaluation criteria (Sec. III.2), and illustrative examples (Sec. III.3) that demonstrate the applicability of machine learning to the cases discussed earlier. Further, in Appendix A we discuss the use a more sophisticated machine learning architecture for better predictability.
II Formulation the of problem
Due to recent theoretical interest in magnetic HHG using spin systems, we will study the input-output relation $h(t)\to Y(t)$ of the incident time-varying magnetic field $h(t)$ and magnetization $Y(t)$ for both the transverse and non-integrable Ising models. Note that time-varying magnetic fields, which are electromagnetic pulses with strong magnetic fields and low electric fields, can now be experimentally produced [74].
@@ II Formulation the of problem
[…]
@@
    with $\text{with }h^{(i)}_{k}=A\sin\left(\omega^{(i)}k\delta t\right),$ (4)
@@
    Modeling the input $\to$ output relation for the non-integrable Ising model (2) based on dataset $\mathcal{D}(A=1.5)$ [Eq. (4)], with the parameters of Hamiltonian (2) set to $h_{1}=0.8\times 0.9045/2$, $h_{z}=0.8\times 0.8090/2$ (see [76]).
@@
    with $\text{with }h^{(i,j)}_{k}=A_{j}\sin\left(\omega^{(i)}k\delta t\right),$
@@ III.1 Model Architecture
The choice of an effective model architecture is important in ML. In our analysis, we use autoencoder TCN architecture. To elucidate our approach, we first explore the encoder-bottleneck-decoder structure inherent to autoencoders within the context of deep neural networks. In Fig. 3, we present an example of a deep neural autoencoder architecture with a configuration of (8-4-2-4-8). This notation means that there are 8 neurons in the input layer followed by 4, 2, and 4 neurons in the subsequent three layers, respectively, and 8 neurons in the output layer. For simplicity, we will denote such a symmetrical architecture in short notation as (8-4-2), which represents the encoder part of the network, with the understanding that the decoder is its mirror image. The encoder takes an input and compresses it into a lower-dimensional representation, called a bottleneck or latent space. The latent space is a smaller vector representation of the input data. The decoder takes this latent representation and reconstructs the output [82, 83, 84]. Similarly, we can have a TCN based autoencoder. An illustration of (8-6-6) autoencoder TCN framework is shown in Fig. 4. We see the temporal 1D convolution layers along with max-pooling layers. In our auto-encoder based architecture, we use kernel/filter$=3$ and dilations ($2,4,8$).
@@ III.3 Illustrations and Discussions
Figure 5: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 1 of Sec. II. Plot (a) shows the $R^{2}$ values vs the frequency $\omega$ for the test data set, i.e., $\{(\omega^{(i)},R^{2}_{i}),\mid i\in\text{TestSet}\}$ [Eqs. (4) and (6)]. The vertical green lines mark the transition frequencies of the time-independent Hamiltonian, demonstrating that the constructed TCN model performs uniformly across the entire frequency range. The purple crosses mark the frequencies $\{\omega^{(i)}\}=\{0.86,1.78,2.45,3.45\}$ of the optical inputs $\mathbf{h}^{(i)}$ [Eq. (4)], labeled as input in subplots (b)-(e). The output time series $\mathbf{Y}^{(i)}$ [Eq. (4)] (labeled as target) induced by the inputs $\mathbf{h}^{(i)}$ and the outputs $\mathbf{\widehat{Y}}^{(i)}$ predicted by the TCN model (pred) are also shown in subplots (b)-(e).
Similarly, in Case 2 of Sec.  II, the minimum stable autoencoder TCN architecture is (12-12-10-10) with 16127 trainable parameters. We again performed a stability assessment by training 10 models of the same architecture and found that for any model, above $85\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ lie above are above $0.90$ (see Fig 6 for one such model). The need for a deeper architecture suggests that the dynamics in Case 2 are more non-linear then in Case 1, resulting in a more intricate output dataset. An independent complexity analysis in the next section will confirm this. Furthermore, we also applied a variational TCN autoencoder (VAE) architecture to case 2 to see if this would provide a significantly better model. It did so and we present the results in appendix A.
Figure 6: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 2 of Sec. II. Plot (a) shows the $R^{2}$ values vs the frequency $\omega$ for the test data set. The vertical green lines mark the transition frequencies of the time-independent Hamiltonian, demonstrating that the constructed TCN model performs uniformly across the entire frequency range. The purple crosses mark the frequencies $\{\omega^{(i)}\}=\{0.86,1.78,2.45,3.45\}$ of the optical inputs $\mathbf{h}^{(i)}$, labeled as input in subplots (b)-(e). The output time series $\mathbf{Y}^{(i)}$ (labeled as target) induced by the inputs $\mathbf{h}^{(i)}$ and the outputs $\mathbf{\widehat{Y}}^{(i)}$ predicted by the TCN model (pred) are also shown in subplots (b)-(e).
@@
Figure 7: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 2 of Sec. II. Plot (a) shows the $R^{2}$ values vs the frequency $\omega$ for the test data set. The vertical green lines mark the transition frequencies of the time-independent Hamiltonian, demonstrating that the constructed TCN model performs uniformly across the entire frequency range. The purple crosses mark the frequencies $\{\omega^{(i)}\}=\{0.86,1.78,2.45,3.45\}$ of the optical inputs $\mathbf{h}^{(i)}$, labeled as input in subplots (b)-(e). The output time series $\mathbf{Y}^{(i)}$ (labeled as target) induced by the inputs $\mathbf{h}^{(i)}$ and the outputs $\mathbf{\widehat{Y}}^{(i)}$ predicted by the TCN model (pred) are also shown in subplots (b)-(e).
@@
Figure 8: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 4 of Sec. II. Plot (a) shows the $R^{2}$ values vs the frequency $\omega$ for the test data set. The vertical green lines mark the transition frequencies of the time-independent Hamiltonian, demonstrating that the constructed TCN model performs uniformly across the entire frequency range. The purple crosses mark the frequencies $\{\omega^{(i)}\}=\{0.86,1.78,2.45,3.45\}$ of the optical inputs $\mathbf{h}^{(i)}$, labeled as input in subplots (b)-(e). The output time series $\mathbf{Y}^{(i)}$ (labeled as target) induced by the inputs $\mathbf{h}^{(i)}$ and the outputs $\mathbf{\widehat{Y}}^{(i)}$ predicted by the TCN model (pred) are also shown in subplots (b)-(e).
@@
In all the cases treated, we ensured a similar train-test split of the data and kept the number of epochs the same. For all our training processes we used the swish activation function [86], the Adam optimizer with a learning rate of 0.001, and and for the loss function we used the Huber loss [87]. Additionally, we employed the technique of changing batch sizes [88]. For all cases we used TCNs with dilation factors of 2, 4, and 8 to effectively capture long-range dependencies in the dataset. This approach allowed us to expand the receptive field of the convolutional layers. Further, in all plots, the input, output, and predicted values are scaled between 0 and 1 because the normalization facilitates a consistent comparison. The plots remain consistent and accurate after applying the inverse transform.
@@
We note that as the complexity of input-output data increases, it is important to examine the behavior of the latent space within the encoder-decoder model. The encoder and decoder collaboratively compress the data into a lower-dimensional representation known as the latent space. When the complexity of the input-output relationship increases, the dimensionality of the latent space must also increase. If the latent dimension is reduced beyond a necessary threshold, the compression quality deteriorates, leading to an insufficient capture of essential features in the latent space. Thus suggests that the size of the smallest latent space that is able to produce an accurate model for the input-output system is a good measure of the complexity of the system. Among the five cases we considered this measure suggests that Case 2 is the most complex. In the next section we evaluate the complexity of the input-output transformations using a more traditional measure of the complexity of time-series. This provides additional evidence that the size of the latent space is a good measure of the system complexity.
III.4 Complexity Analysis

### v2

@@ I Introduction
Optics studies light and its interactions with matter. As depicted in Fig 1, the study comprises three main components: a quantum system, the incident pulse of light (optical input), and the system’s response (optical output) to the incident pulse. The system’s response usually varies depending on the input signal. The type of input signal can determine whether the dynamical system behaves linearly, non-linearly, or chaotically (a specific type of non-linear behavior characterized by extreme sensitivity) [1]. If the input is weak, we have what is called linear optics. If the input intensity increases, we have a non-linear response/output signal, and the output can be expressed as a perturbation series expansion of the input [2]. As the intensity of the input field increases further, the system exhibits even more pronounced non-linearities requiring non-pertrubative treatment [3], one of the most significant being High Harmonic Generation (HHG). This strong-field laser–matter interaction resulting in HHG has attracted a lot of attention, and its discovery has been recognized with the 2023 Nobel Prize in Physics. HHG has been extensively studied in various media, including gases, semiconductors, nano-structures, metal, strongly correlated materials, and solids [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 10, 17, 18, 19, 20, 21, 22, 23, 24]. More recently, there is mounting interest in HHG from quantum spin systems due to their potential applications in probing spin dynamics and developing new laser sources in the THz regime [25, 26, 27, 28].
The optical output (Fig 1) is obtained theoretically by solving the time-dependent Schrödinger equation. Solving this equation rapidly becomes prohibitively expensive, as finding its solution is exponentially difficult with respect to the size of the quantum system with which the input pulse interacts. The complexity increases further if the quantum system is undergoing a phase transition.
@@
Light-matter interactions during phase transitions have been observed experimentally [29]. Understanding phase transitions is important for various reasons [30, 31, 32, 33]. For example, in quantum computing, the transverse field Ising model is significant for quantum annealing, where phase transitions can influence the performance and efficiency of quantum algorithms [34]. The transverse Ising model is well known to exhibit a quantum phase transition and is analytically solvable [35]. However, an Ising model with mixed transverse and longitudinal magnetic fields can only be solved numerically  [36]. In such regimes, small changes in system parameters can lead to large-scale changes in the system’s dynamics, requiring extremely precise calculations that are challenging to achieve by numerically solving the Schrödinger equation. Alternative methods are needed to accurately model the behavior of such systems.
@@
In this paper, we propose a novel data-driven methodology to model the input $\to$ output dynamics of optical systems using machine learning. We utilize Temporal Convolutional Networks (TCNs) to capture the intricate interactions in many-body quantum systems by formulating optics as an input-output time series problem. While the presented methodology is generally applicable across different phenomena, we demonstrate our approach using the quantum Ising model as a test case. Additionally, we show how distinct physical regimes, including phase transitions, can be captured with machine learning models (see Fig. 1), tailored to the complexity of each regime.
@@
A conventional application of machine learning in time-series analysis is forecasting [39, 44]. A notable example is Ref. [45], where using a historical record (e.g., the training data shown in Fig. 2) for a time-domain output $Y(t)$, a model was constructed to predict $Y(t)$ for future times (see Fig. 2). It is important to note that this modeling does not utilize information across different input pulses. Hence, each time the incident laser pulse is changed, a new model needs to be trained. In the present work, we depart from this forecasting approach by training a TCN to learn the entire input $\to$ output relationship, employing a dataset of pairs of input and output pulses.
Figure 2: Schematic picture of time series modelling using TCN/LSTM: $Y(t)$ is the optical output as a function of time. Atomic units (a.u.) are used throughout. The blue portion depicts the historical data that serves as the training data for the model, the red portion depicts the test data or the unseen data for the model and the green curve depicts the model’s predicted data whose closeness to the test data is a measure of the performance of the model.
@@
The application of machine learning to light-matter interactions has been prolific. However, unlike our current work, the previous studies have not enforced causality. These include applications in harmonic generation microscopy [57], nonlinear spectroscopy in solids [58], predicting characteristics of harmonic generation in plasmas using particle-in-cell simulations [59], predicting HHG emission from molecules and inversely predicting molecular parameters from the corresponding spectra [60], time series forecasting of nonlinear spectra [45], and predicting HHG from the spatially structured input fields [61]. Additionally, deep neural networks have been employed in reconstruction of ultrashort pulses [62, 63], prediction of pulse properties generated by a free-electron laser [64], and denoising measured photoelectron spectra [65].
@@
The rest of the paper is organized as follows: In Sec. II, we comprehensively formulate the problem under investigation and provide details of the Ising system under consideration and various cases associated with it. Section III is dedicated to presenting our results; we state details about the machine learning model architectures employed (Sec. III.1), evaluation criteria (Sec. III.2), illustrative examples (Sec. III.3) that demonstrate the applicability of machine learning to the cases discussed earlier. In Sec. III.4, we discuss the use of a more sophisticated generative machine learning architecture (Variational Autoencoder TCN) for better predictability. Further, we discuss in Sec. III.5 how the latent space of the architecture can be used to understand the complexity of the data.
II Formulation of the problem
Due to recent theoretical and experimental interest in magnetic HHG using spin systems, we will study the input-output relation $h(t)\to Y(t)$ of the incident time-varying magnetic field $h(t)$ and magnetization $Y(t)$ for both the transverse and non-integrable Ising models. Note that time-varying magnetic fields, which are electromagnetic pulses with strong magnetic fields and low electric fields, can now be experimentally produced [70].
@@ II Formulation of the problem
where $\left\langle\quad\right\rangle$ denotes the averaging over the time-dependent wave function.
@@
    $\text{with }h^{(i)}_{k}=A\sin\left(\omega^{(i)}k\delta t\right),$ (4)
@@
    Modeling the input $\to$ output relation for the non-integrable Ising model (2) based on dataset $\mathcal{D}(A=1.5)$ [Eq. (4)], with the parameters of Hamiltonian (2) set to $h_{1}=-0.8\times 1.05/2$, $h_{z}=0.8\times 0.5/2$ (see [72]).
@@
    $\text{with }h^{(i,j)}_{k}=A_{j}\sin\left(\omega^{(i)}k\delta t\right),$
@@ III Results
The complete set of codes that reproduce all results presented below is available at [78].
@@ III.1 Model Architecture
The choice of an effective model architecture is important in ML. In our analysis, we use autoencoder TCN architecture. To elucidate our approach, we first explore the encoder-bottleneck-decoder structure inherent to autoencoders within the context of deep neural networks. In Fig. 3, we present an example of a deep neural autoencoder architecture with a configuration of (8-4-2-4-8). This notation means that there are 8 neurons in the input layer followed by 4, 2, and 4 neurons in the subsequent three layers, respectively, and 8 neurons in the output layer. For simplicity, we will denote such a symmetrical architecture in short notation as (8-4-2), which represents the encoder part of the network, with the understanding that the decoder is its mirror image. The encoder takes an input and compresses it into a lower-dimensional representation, called a bottleneck or latent space. The latent space is a smaller vector representation of the input data. The decoder takes this latent representation and reconstructs the output [79, 80, 81]. Similarly, we can have a TCN based autoencoder. An illustration of (8-6-6) autoencoder TCN framework is shown in Fig. 4. We see the temporal 1D convolution layers along with max-pooling layers. In our auto-encoder based architecture, we use kernel/filter=3=3 and dilations ($2,4,8$).
@@ III.3 Illustrations and Discussions
In Fig. 5(a), we observe an oscillatory behavior for $R^{2}$ at low frequencies $\omega$. To determine whether this oscillation reveals any underlying physics, we compare it with the characteristic transition frequencies of the quantum system $\omega_{kl}=(E_{k}-E_{l})/\hbar$, $E_{k}>E_{l}$, (recall that $\hbar=1$ in atomic units), where $E_{m}$ are eigenenergies (i.e., eigenvalues) of the time-independent part of the Hamiltonian. Physically, $\omega_{kl}$ denotes the energies of photons required to drive transitions between energy levels $E_{l}$ and $E_{k}$. In Fig. 5(a), the vertical green lines mark the transition frequencies $\omega_{kl}$, which cluster into bands. Since no correlation is observed between $R^{2}$ and the transition frequencies, we conclude that the time-series model performs uniformly across the entire frequency range.
Figure 5: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 1 of Sec. II. Plot (a) shows the $R^{2}$ values vs the frequency $\omega$ for the test data set, i.e., $\{(\omega^{(i)},R^{2}_{i}),\mid i\in\text{TestSet}\}$ [Eqs. (4) and (6)]. The vertical green lines mark the transition frequencies of the time-independent Hamiltonian, demonstrating that the constructed time series model performs uniformly across the entire frequency range. The purple crosses mark the frequencies $\{\omega^{(i)}\}=\{0.86,1.78,2.45,3.45\}$ of the optical inputs $\mathbf{h}^{(i)}$ [Eq. (4)], labeled as input in subplots (b)-(e). The output time series $\mathbf{Y}^{(i)}$ [Eq. (4)] (labeled as target) induced by the inputs $\mathbf{h}^{(i)}$ and the outputs $\mathbf{\widehat{Y}}^{(i)}$ predicted by the time series model (pred) are also shown in subplots (b)-(e).
Similarly, in Case 2 of Sec.  II, the minimum stable autoencoder TCN architecture is (12-12-10-10) with 16127 trainable parameters. We again performed a stability assessment by training 10 models of the same architecture and found that for any model, above $85\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ lie above are above $0.90$ (see Fig 6 for one such model). The need for a deeper architecture suggests that the dynamics in Case 2 are more non-linear then in Case 1, resulting in a more intricate output dataset. An independent complexity analysis in the next section will confirm this. Furthermore, we also applied a variational TCN autoencoder (VAE) architecture to case 2 to see if this would provide a significantly better model. It did so and we present the results in in III.4.
Figure 6: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 2 of Sec. II. See the caption of Fig. 5 for further details.
@@
Figure 7: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 3 of Sec. II. See the caption of Fig. 5 for further details.
@@
Figure 8: Performance illustration of a temporal convolutional network (TCN) model constructed for Case 4 of Sec. II. See the caption of Fig. 5 for further details.
@@
In all the cases treated, we ensured a similar train-test split of the data and kept the number of epochs the same. For all our training processes we used the swish activation function [83], the Adam optimizer with a learning rate of 0.001, and for the loss function we used the Huber loss [84]. Additionally, we employed the technique of changing batch sizes [85]. For all cases we used TCNs with dilation factors of 2, 4, and 8 to effectively capture long-range dependencies in the dataset. This approach allowed us to expand the receptive field of the convolutional layers. Further, in all plots, the input, output, and predicted values are scaled between 0 and 1 because the normalization facilitates a consistent comparison. The plots remain consistent and accurate after applying the inverse transform.
We have summarized the training details and the results in Tables 1 and 2, respectively.
Model Type Temporal Convolutional Networks
Activation Function Swish
Loss Function Huber Loss ([Reference 87])
Learning Rate 0.001
Optimizer Adam
Train-Test Split Similar split across all cases
Epochs 250
Dilation Factors 2, 4, and 8 (used to capture long-range dependencies)
Table 1: Summary of Training Configurations
@@
We note that as the complexity of input-output data increases, it is important to examine the behavior of the latent space within the encoder-decoder model. The encoder and decoder collaboratively compress the data into a lower-dimensional representation known as the latent space. When the complexity of the input-output relationship increases, the dimensionality of the latent space must also increase. If the latent dimension is reduced beyond a necessary threshold, the compression quality deteriorates, leading to an insufficient capture of essential features in the latent space. Thus suggests that the size of the smallest latent space that is able to produce an accurate model for the input-output system is a good measure of the complexity of the system. Among the five cases we considered this measure suggests that Case 2 is the most complex. In Sec. III.5, we evaluate the complexity of the input-output transformations using a more traditional measure of the complexity of time-series. This provides additional evidence that the size of the latent space is a good measure of the system complexity.
Cases Hamiltonian Amplitude Algorithm Architecture Parameters Loss Function Epochs Result
1 Eq. (1) $A=1$ TCN 5-5-3 2082 Huber Loss 250 over $90\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.98$
2 Eq. (1) $A=10$ TCN 12-12-10-10 16127 Huber Loss 250 over $85\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.90$
2 Eq. (1) $A=10$ VAE-TCN 12-12-10-10 34127 Huber Loss 250 over $90\%-95\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.90$
3 Eq. (2) $A=1.5$ TCN 5-5-4 2234 Huber Loss 250 over $95\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.95$
4 Eq. (2) $A=2.5$ TCN 8-8-6 5439 Huber Loss 250 over $90\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.95$
5 Eq. (2) $A_{1}=1$ $A_{m}=1.7$ TCN 8-8-4 3221 Huber Loss 250 over $88\%$ of $\{R^{2}_{i}|i\in\text{ TestSet}\}$ consistently exceed $0.95$
Table 2: Performance evaluation of time series algorithms for the cases listed in Sec. II, studying different driven Ising systems. The results highlight the proportion of test set predictions achieving high R-squared values.
III.4 Variational TCN Autoencoder
Figure 9: Performance illustration of a variational temporal convolutional network autoencoder (VAE TCN) model constructed for Case 2 of Sec. II. See the caption of Fig. 5 for further details.
III.5 Complexity Analysis